given: 5x+4=39
prove: 2x-3=11
what the statements and reason
1
Expert's answer
2012-09-04T09:47:58-0400
5x+4=39 let's find x from here
5x=39-4=35
x=7
Now we can input the value we've calculated for x into the second equation:
2x-3=11
input x=7:& 2*7-3=14-3=11 and that's it